OpenStudy (anonymous):

hi.choose a number X at random from the interval [0,2].What is the probability that i)the first digit to the right of the decimal is 3? ii)the second digit is a 7?

OpenStudy (anonymous):

i believe it should be \(\frac{1}{10}\) for both

OpenStudy (anonymous):

how do you get to that?

OpenStudy (apoorvk):

the digits can be any one from 1,2,3,4,5....,9,0. So, 1/10.

OpenStudy (unklerhaukus):

what are the possibilities for the number X?

OpenStudy (unklerhaukus):

is X an integer \(\mathbb Z\)? or any real number \(\mathbb R?\)

OpenStudy (anonymous):

i think it is an integer

OpenStudy (unklerhaukus):

oh it is a decimal \(\mathbb D \). the possibilities for \(X\in[0,2]\) are \(\{0.0,0.1,0.2,0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9,1.0,\cdots\)\[ \quad\qquad \cdots 1.1, 1.2, 1.3, 1.4, 1.5, 1.6, 1.7, 1.8, 1.9, 2.0\}\] there are \(\textbf{two}\) occurrences of each \[\{ 1, 2, 3, 4, 5, 6, 7, 8, 9, \}\]and \(\textbf {three}\) occurrences of \(\{0,\}\) as the first digit after the decimal place this set has ten possible outcomes but not each is equally likely

OpenStudy (unklerhaukus):

the probability of a three as the first place after the decimal is the number of occurrence of three divided by all the possible occurrences \[P(3)=\frac{2}{21}\]

OpenStudy (anonymous):

ok.don't get it entirely bt i'll digest it.thanx @UnkleRhaukus ,@apoorvk ,@satellite73

OpenStudy (anonymous):

thanx @apoorvk @satellite73

OpenStudy (apoorvk):

Actually it was @UnkleRhaukus really who did a great job here, as he always does!

OpenStudy (apoorvk):

No worries still ;)

OpenStudy (unklerhaukus):

im not entirely sure X is a decimal

OpenStudy (unklerhaukus):

i think for ii) ' the second digit ' is again referring to the first to the right digit after the decimal place,

OpenStudy (apoorvk):

Well, it asks for the 'digit' not what the no. could be. and a digit could only be out of 10 possible options. WAIT. If the last digit after the decimal is '0', then it would not be significant right?

OpenStudy (unklerhaukus):

hmmm, that is a good question

OpenStudy (anonymous):

thank u all

OpenStudy (anonymous):

guess its significance is that u can choose a whole 0 meaning there is no digit after the decimal place.jst like u can choose a whole 1 and 2

OpenStudy (anonymous):

maybe i misunderstood the question, but i thought it asked "pick a number between 0 and 2 at random"

OpenStudy (anonymous):

not a number with only two decimal places, but any number between 0 and 2 for example \(\frac{1}{e}\) or \(\frac{\pi}{6}\)

OpenStudy (anonymous):

in which case the probability that the number will have a 3 to the right of the decimal place means it begins \(0.3...\) or \(1.3...\) and the digits will go on forever. one tenth will begin \(0.0...\) or \(1.0...\) and one tenth will begin \(0.1...\) or \(1.1...\) and so on. therefore if i am interpreting the question correctly, one tenth will begin \(0.3...\) or \(1.3...\)

OpenStudy (anonymous):

@UnkleRhaukus i think there might be an error in your calculation as follows. since the number is chosen in the interval \([0,2]\) a continuum, this is a length problem. length of the interval divided by total length. the probability that the number 2 is chosen is zero

OpenStudy (anonymous):

i could be wrong

OpenStudy (unklerhaukus):

if 2 were excluded wouldn't the interval be \([0,2)\)

OpenStudy (anonymous):

the problem is to pick a number at random in the interval [0,2] which is the same as the interval (0,2) as far as probability is concerned. there are an uncountable infinite number of numbers in that interval of length 2, and the probability that any specific number is chosen is zero

OpenStudy (anonymous):

for example, if i say "pick a number in the interval [0,2] what is the probability that it is between \(\frac{1}{2}\) and 1, then the solution is the length of the favorable interval, namely \(\frac{1}{4}\) divided by the total length, so the solution is \(\frac{1}{8}\) the inclusion or exclusion of the endpoints in not important

OpenStudy (unklerhaukus):

1/4?

OpenStudy (anonymous):

ok that was wrong, i meant \(\frac{1}{2}\) and probability is \(\frac{1}{4}\) a mistake on my part, but idea is clear right?

OpenStudy (anonymous):

in other words, on the interval \([0,2]\) \(P(1.3<x<1.4)=P(1.3\leq x\leq 1.4)\)

OpenStudy (unklerhaukus):

then why were the end points include in the interval \([0,2]\) instead of \((0,2)\), or do they mean the same thing in a continuous random distribution?

OpenStudy (anonymous):

for a continuous probability it doesn't make any difference.

OpenStudy (anonymous):

the probability any specific number is chosen is zero, the solution is the measure of the length of the favorable interval divided by the length of the total interval. the measure of (0, 2) and [0,2] are the same, both intervals have length 2

OpenStudy (anonymous):

in fact, since the rational numbers have measure zero, you could not only exclude the end point, but could exclude all rational numbers and the answer would not change

OpenStudy (unklerhaukus):

well for my working above i have assumed discrete probability , \(X\in \mathbb D\)

OpenStudy (unklerhaukus):

i now am becoming convinced that this was a mistake

OpenStudy (anonymous):

if the only possible numbers chosen were \[\{0.0,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\] \[,1.0,1.1,1.2,1.3,1.4,1.5,1.6,1.7,1.8,1.9,2\}\] then you would be correct for sure

OpenStudy (anonymous):

you have 21 discrete numbers, of which 2 have the first digit after the decimal a 3 but i am fairly sure that when it says "pick a number at random" number mean s real number.

OpenStudy (unklerhaukus):

ok i concede \( X\in \mathbb R\) \[P(3 \text{ is the first digit to the right of the decomal place})=\frac1{10}\]